x+1/x=5, then find the value of x^16+1/x^16 ?


1
Expert's answer
2021-01-29T05:15:56-0500

We have given with

"x+1\/x=5"

So, "(x+1\/x)^2=(5)^2"

Expanding the bracket by using formula

"(a+b)^2=a^2+2ab+b^2"

We get ,

"x^2+2(x)(1\/x)+(1\/x^2)=25"

Simplifying we have,

"x^2+1\/x^2=23"

Taking square on both sides we get

"(x^2+1\/x^2)^2=(23 )^2"

"=x^4+2(x^4)(1\/x^4)+(1\/x^4)=529" ,

"x^4+1\/x^4=529-2"

"=x^4+1\/x^4=527"

Again Taking square on both sides we get

"(x^4+1\/x^4)^2=(527)^2"

"=x^8+2(x^8)(1\/x^8)+1\/x^8=277729" ,

"x^8+1\/x^8=277729-2"

"=x^8+1\/x^8=277727"

Again ,Taking square on both sides,we have

"x^{16}+1\/x^{16}=(277727)^2-2" So,
Answer:

"(277727)^2-2"